# HOPSE: Scalable Higher-Order Positional and Structural Encoder for Combinatorial Representations

### v1

@@ A.2 Neighborhood functions and augmented Hasse graphs
Given a combinatorial complex ${\mathcal{T}}=(\mathcal{V},\mathcal{C},{\textnormal{k}})$, a neighborhood $\mathcal{N}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{C})$ on a combinatorial complex is a function that assigns to each cell $\sigma\in\mathcal{C}$ a collection of neighbor cells $\mathcal{N}(\sigma)\subset\mathcal{C}\cup\emptyset$.
Each neighborhood can be used to create a Strictly Augmented Hasse graph.
@@
Each particular neighborhood $\mathcal{N}$ induces a strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}=(\mathcal{C}_{\mathcal{N}},\mathcal{E}_{\mathcal{N}})$ on $\mathcal{T}$ [39, 23], defined as the directed graph whose nodes and edges are given, respectively, by $\mathcal{C}_{\mathcal{N}}=\{\sigma\in\mathcal{C}\,|\,\mathcal{N}(\sigma)\neq\emptyset\}$ and $\mathcal{E}_{\mathcal{N}}=\{(\tau,\sigma)\,|\,\tau\in\mathcal{N}(\sigma)\}$.
Examples of neighborhood functions are incidences (connecting cells with different ranks) and adjacencies (connecting cells with the same rank), although other neighborhoods can be defined for specific tasks [3]. More in detail, given a combinatorial complex ${\mathcal{T}}$:
@@
$\begin{split}\mathcal{I}_{s\to t}(\sigma)=\left\{\begin{array}[]{ll}\{\tau\in\mathcal{C}\,|\,\textrm{rk}(\tau)=s,\sigma\subset\tau\}&\text{ if $t<s$;}\\ \{\tau\in\mathcal{C}\,|\,\textrm{rk}(\tau)=s,\tau\subset\sigma\}&\text{ if $t>s$.}\end{array}\right.\end{split}$ (7)
@@
Appendix B Counting neighborhoods and message passing routes
@@
B.1 Neighborhood Types in Combinatorial Complexes
Combinatorial representations allow for a variety of neighborhood types, depending on how cells of different ranks interact. For a 2-dimensional combinatorial complex (with 00-, $1$-, and $2$-cells), Figure 3 illustrates all possible neighborhood types. These relations can be grouped into two main groups:
@@ B.1 Neighborhood Types in Combinatorial Complexes
B.2 Counting neighborhoods
The combinatorial complex depicted in Figure 3 has rank $R=\dim({\mathcal{T}})=2$, which results in 12 distinct neighborhood types. These neighborhoods can be divided into the two categories from above: (i) incidence relations and (ii) adjacency relations.
@@ B.2 Counting neighborhoods
distinct neighborhood types. For the example in Figure 3, where $R=2$, this formula recovers the 12 neighborhoods observed.
B.3 Message passing routes
The number of neighborhoods increases quadratically with the maximum rank, which may initially seem manageable, especially since most practical datasets involve low-rank structures (e.g., $R=2$ or $R=3$). However, the true combinatorial complexity arises when considering how these neighborhoods can be composed. To formalize this notion, the following definition is introduced.
Definition B.1 (Message passing route).
A message passing route is an ordered set of neighborhoods that a model uses to propagate information across a data structure.
Information may flow intra-rank, through adjacency relations, or inter-rank, through incidence relations. For example, a network might use the message passing route $[{\mathcal{I}}_{2\to 1},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{0\to 2},{\mathcal{A}}_{0,1}]$ to transmit information from faces to edges, then from edges to nodes, from nodes back to faces, and finally between nodes connected by an edge.
@@ B.3 Message passing routes
Definition B.2 (Minimal message passing route).
@@
Minimal routes are of particular interest, as they guarantee full interdependence between all cells—a desirable property for expressive models. Moreover, minimal length implies simplicity in design, yielding the most compact routes that preserve this full dependency. For instance, the previously introduced route $[{\mathcal{I}}_{2\to 1},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{0\to 2},{\mathcal{A}}_{0,1}]$ can be transformed into a minimal message passing route by removing the adjacency ${\mathcal{A}}_{0,1}$. In this reduced route, information from all cells becomes accessible to all others within two message passing steps. In fact, consider the case of 00-cells: after the first step, they depend on $1$-cells; after the second step, since $1$-cells now incorporate information from $2$-cells, the 00-cells become indirectly connected to them as well.
By definition, minimal message passing routes rely exclusively on incidence relations, as adjacencies do not propagate information inter-rank. Since each incidence targets a specific rank, and every rank must be updated once, minimal routes necessarily have length $R+1$. To enumerate all possible minimal routes, consider the set of ranks $T=\{0,1,\dots,R\}$, of size $R+1$. There are $(R+1)!$ distinct permutations of these ranks, each defining a unique order in which the ranks are updated. Given such a permutation $\{r_{0},r_{1},\dots,r_{R}\}$, a corresponding minimal message passing route can be constructed by selecting the sequence of incidence relations ${\mathcal{I}}_{r_{j-1}\to r_{j}}$ for $j=1,\dots,R$, and closing the cycle with ${\mathcal{I}}_{r_{R}\to r_{0}}$. This construction yields exactly one valid minimal route per permutation, resulting in a total of $N=(R+1)!$ such routes. In the case of graphs, where $R=1$, this results in exactly two minimal message passing routes—an intuitively expected outcome.
B.4 Beyond minimal routes: including adjacencies
Minimal message passing routes represent only a small subset of the possible routes that can be constructed using the full set of neighborhoods in a combinatorial complex. A natural extension involves augmenting these minimal routes with adjacency operations—a common practice in many TDL architectures. For simplicity, the analysis here is restricted to the case where adjacencies are applied strictly after all incidence relations.

### v2

@@ A.2 Neighborhood functions and augmented Hasse graphs
Given a combinatorial complex ${\mathcal{T}}=(\mathcal{V},\mathcal{C},{\textnormal{k}})$, a neighborhood function $\mathcal{N}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{C})$ on a combinatorial complex is a function that assigns to each cell $\sigma\in\mathcal{C}$ a collection of neighbor cells $\mathcal{N}(\sigma)\subset\mathcal{C}\cup\emptyset$.
Each neighborhood function can be used to create a Strictly Augmented Hasse graph.
@@
Each particular neighborhood function $\mathcal{N}$ induces a strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}=(\mathcal{C}_{\mathcal{N}},\mathcal{E}_{\mathcal{N}})$ on $\mathcal{T}$ [Papillon et al., 2024, Hajij et al., 2023], defined as the directed graph whose nodes and edges are given, respectively, by $\mathcal{C}_{\mathcal{N}}=\{\sigma\in\mathcal{C}\,|\,\mathcal{N}(\sigma)\neq\emptyset\}$ and $\mathcal{E}_{\mathcal{N}}=\{(\tau,\sigma)\,|\,\tau\in\mathcal{N}(\sigma)\}$.
Examples of neighborhood functions are incidences (connecting cells with different ranks) and adjacencies (connecting cells with the same rank), although other neighborhood functions can be defined for specific tasks [Battiloro et al., 2025]. More in detail, given a combinatorial complex $\mathcal{T}$}:
@@
[…]
@@
A.3 Morphisms between CCs
Before defining isomorphism between CCs, the definition of homomorphism between CCs is needed
Definition A.12 (CC-Homomorphism (Definition 10 Hajij et al. [2023]).
A homomorphism from a $CC(S_{1},\mathcal{X}_{1},\mathrm{rk}_{1})$ to another $CC(S_{2},\mathcal{X}_{2},\mathrm{rk}_{2})$ is a function $f:\mathcal{X}_{1}\to\mathcal{X}_{2}$ that sasfies the following conditions:
  1. 1.
    If $x,y\in\mathcal{X}_{1}$ satisfy $x\subseteq y$, then $f(x)\subseteq f(y)$
  2. 2.
    If $x\in\mathcal{X}_{1}$, then $\mathrm{rk}_{1}(x)\geq\mathrm{rk}_{2}(f(x))$.
However, it is not enough to work in the attributed setting. A CC with attributes assigned to cells is
Definition A.13 (Attributed CC [Papillon et al., 2024]).
An attributed combinatorial complex $({\mathcal{C}},\ell)$ is a CC denoted $\mathcal{C}$} equipped with a cell coloring function $\ell:{\mathcal{C}}\to\Sigma$ where $\Sigma$ is an arbitrary normed space. For a cell $\sigma$ that is part of $\mathcal{C}$}, $\ell(\sigma)$ denotes the label or color of $\sigma$.
Now note that there also exist homomorphisms induced by the neighborhood functions that can be applied to the CCs. The definition of such is
Definition A.14 (CC-Homomorphism induced by $(\mathcal{N}_{1},\mathcal{N}_{2})$ [Papillon et al., 2024]).
A homomorphism from a $CC$ denoted $\mathcal{CC}_{1}=(\mathcal{V}_{1},\mathcal{C}_{1},\mathrm{rk}_{1})$ with a neighborhood function $\mathcal{N}_{1}$ to another $CC$ denoted $\mathcal{CC}_{2}=(\mathcal{V}_{2},\mathcal{C}_{2},\mathrm{rk}_{2})$ with a neighborhood function $\mathcal{N}_{2}$ is a function $f:\mathcal{C}_{1}\to\mathcal{C}_{2}$ that satisfies the following condition: If $\sigma,\tau\in\mathcal{C}_{1}$ and $\tau\in\mathcal{N}_{1}(\sigma)$ then $f(\tau)\in\mathcal{N}_{2}(f(\sigma))$.
This definition extends to attributed CCs with the following condition: if $\sigma,\tau\in\mathcal{C}_{1}$ have the same label then their image in $\mathcal{C}_{2}$ also has the same label, i.e for $\ell:\mathcal{C}_{1}\to\Sigma$ and $\ell^{\prime}:\mathcal{C}_{2}\to\Sigma$, if $\ell(\sigma)=\ell(\tau)$ then $f(\ell^{\prime}(\sigma))=f(\ell^{\prime}(\tau))$. The neighborhood functions also induce Hasse graphs. In fact, homomorphisms between CCs induced by neighborhood functions are one-to-one to homomorphisms of the Hasse graphs induced by the same neighborhood functions.
Proposition A.15 (Proposition B.4 in [Papillon et al., 2024]).
For every CC-Homomorphism $f$ from $\mathcal{C}_{1}$ to $\mathcal{C}_{2}$ induced by the neighborhood functions $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, with domains $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ respectively, there exists a unique homomorphism between the strictly augmented Hasse graphs $\mathcal{G}_{\mathcal{N}_{1}}$ and $\mathcal{G}_{\mathcal{N}_{2}}$.
The generality of combinatorial complexes complicates their direct isomorphic comparison. Interestingly, they can be decomposed and evaluated on their constituting pieces, Hasse graphs. Thus, isomorphisms can be induced by neighborhood functions, which have a one-to-one correspondance with Hasse graphs, which give an equivalence relationship between CCs. This type of equivalence relationship is definited as
Definition A.16 (CC-Isomorphism induced by $(\mathcal{N}_{1},\mathcal{N}_{2})$ [Papillon et al., 2024]).
A isomorphism from a $CC$ denoted $\mathcal{CC}_{1}=(\mathcal{V}_{1},\mathcal{C}_{1},\mathrm{rk}_{1})$ with a neighborhood function $\mathcal{N}_{1}$ to another $CC$ denoted $\mathcal{CC}_{2}=(\mathcal{V}_{2},\mathcal{C}_{2},\mathrm{rk}_{2})$ with a neighborhood function $\mathcal{N}_{2}$ is an invertible CC-Homomorphism induced by $(\mathcal{N}_{1},\mathcal{N}_{2})$ whose inverse is a CC-Isomorphism induced by $(\mathcal{N}_{2},\mathcal{N}_{1})$. In the labeled CC case, then the map also preserves labels.
Definition A.17 (CC-Isomorphism).
A isomorphism from a $CC$ denoted $\mathcal{CC}_{1}=(\mathcal{V}_{1},\mathcal{C}_{1},\mathrm{rk}_{1})$ to another $CC$ denoted $\mathcal{CC}_{2}=(\mathcal{V}_{2},\mathcal{C}_{2},\mathrm{rk}_{2})$ is a bijection $f$ that fulfils:
  • $\forall\sigma,\tau\in\mathcal{C}_{1}$, if $\sigma\subseteq\tau$ then $f(\sigma)\subseteq f(\tau)$,
  • $\forall\sigma\in\mathcal{C}_{1}$ it holds that $\mathrm{rk}_{1}(\sigma)=\mathrm{rk}_{2}(f(\sigma))$.
If there exists such an $f$, the CCs are deemed isomorphic, denoted $CC_{1}\simeq CC_{2}$.
A.4 Characterization and invariants
One of the ways to determine if graphs are different is thorugh graph invariants. These are functions that evaluate to the same output if the graphs are the same under an equivalence relationship, i.e. isomorphism.
Definition A.18 (Graph invariant).
A graph invariant is a function $\xi$ on a graph that assigns the same value to isomorphic graphs. For a fixed $G$ and any other $G^{\prime}$ such that $G\simeq G^{\prime}$ we have $\xi(G)=\xi(G^{\prime})$. A graph invariant distinguishes non-isomorphic graphs $G$ and $G^{\prime}$, denoted $G\not\equiv_{\xi}G^{\prime}$ if $\xi(G)\neq\xi(G^{\prime})$ and conversely if $G$ and $G^{\prime}$ are isomorphic then $G\equiv_{\xi}G^{\prime}$ if $\xi(G)=\xi(G^{\prime})$.
An equivalent concept can be formulated in terms of combinatorial complexes, as we have a notion of equivalence from Definition˜A.17.
Definition A.19 (CC-Invariant).
A CC-Invariant is a function $\xi$ on a CC that assigns the same value to isomorphic CCs. For a fixed combinatorial complex $CC_{1}$ and any other $CC_{2}$ such that $CC_{1}\simeq CC_{2}$ we have $\xi(CC_{1})=\xi(CC_{2}^{\prime})$. As with graph invariants, this function helps distinguish non-isomorphic CCs.
Now notice that a CC-Invariant induced by a neighborhood tuple $(\mathcal{N}_{1},\mathcal{N}_{2})$ is simply a graph invariant on the resuling Hasse induced by the neighborhoods. With this in mind, Hasse graphs can be used as piece-wise invariants to compose an invariant over the whole complex.
Appendix B Theoretical motivation
Proposition B.1.
Let $\mathcal{CC}_{1}=(\mathcal{V}_{1},\mathcal{C}_{1},\mathrm{rk}_{1})$ and $\mathcal{CC}_{2}=(\mathcal{V}_{2},\mathcal{C}_{2},\mathrm{rk}_{2})$. Let $\mathcal{N}_{1}$ be a neighborhood function on $\mathcal{CC}_{1}$ and $\mathcal{N}_{2}$ be a neighborhood function on $\mathcal{CC}_{2}$. For every CC-Isomorphism $f:\mathcal{C}_{1}\to\mathcal{C}_{2}$ induced by $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, there exists a unique graph isomorphism between their respective strictly augmented Hasse graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$.
Proof.
Note that this is trivial since every isomorphism is also a homomorphism and by Proposition˜A.15 it is unique. ∎
First, we denote two special kinds of neighborhood functions. An incidence neighborhood function $\mathcal{N}$ is such that $\forall\sigma,\tau\in\mathcal{C}$, where $\sigma\in\mathcal{N}(\tau)$ or $\tau\in\mathcal{N}(\sigma)$ only one of these conditions is true:
  1. 1.
    $\mathrm{rk}_{1}(\sigma)>\mathrm{rk}_{1}(\tau)$,
  2. 2.
    $\mathrm{rk}_{1}(\sigma)<\mathrm{rk}_{1}(\tau)$.
Conversly, a adjacency neighborhood function $\mathcal{N}$ is such that $\forall\sigma,\tau\in\mathcal{C}$, where $\sigma\in\mathcal{N}(\tau)$ or $\tau\in\mathcal{N}(\sigma)$ then $\mathrm{rk}_{1}(\sigma)=\mathrm{rk}_{1}(\tau)$.
Proposition B.2.
Let $CC_{1}$ and $CC_{2}$ be two CCs where $CC_{1}\not\simeq CC_{2}$. There exists a pair of neighborhood functions $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, on $CC_{1}$ and $CC_{2}$ respectively, such that the induced Hasse graphs $\mathcal{G}_{\mathcal{N}_{1}}$ and $\mathcal{G}_{\mathcal{N}_{2}}$ are non-isomorphic.
Proof.
Assume that for any choice of functions $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ then $\mathcal{G}_{\mathcal{N}_{1}}$ and $\mathcal{G}_{\mathcal{N}_{2}}$ are isomorphic. This means that there is an isomorphism $\hat{f}:\mathcal{G}_{\mathcal{N}_{1}}\to\mathcal{G}_{\mathcal{N}_{2}}$. This means that
$\forall u,v\in V(\mathcal{G}_{\mathcal{N}_{1}})\text{ s.t. }(u,v)\in E(\mathcal{G}_{\mathcal{N}_{1}})\implies(\hat{f}(u),\hat{f}(v))\in E(\mathcal{G}_{\mathcal{N}_{2}})$ (9)
and
$\forall u^{\prime},v^{\prime}\in V(\mathcal{G}_{\mathcal{N}_{2}})\text{ s.t. }(u^{\prime},v^{\prime})\in E(\mathcal{G}_{\mathcal{N}_{2}})\implies(\hat{f}^{-1}(u),\hat{f}^{-1}(v^{\prime}))\in E(\mathcal{G}_{\mathcal{N}_{1}}).$ (10)
Note that each isomorphism between Hasse graph also uniquely defines a CC-Isomorphism $f$ induced by the neighborhood functions $(\mathcal{N}_{1},\mathcal{N}_{2})$. Now, notice that only for a specific family of neighborhood functions, this maps encode rank information. This family are the incidence neighborhood functions of the form $\mathcal{I}_{s\to t}$ where $s\neq t$. Since $\mathcal{G}_{\mathcal{N}_{1}}$ and $\mathcal{G}_{\mathcal{N}_{2}}$ are directed graphs that means that rank is presserved in $f$. In other words, this implies that there is a CC-Isomorphism $g$ since for any restriction over the domain $\mathcal{C}_{1}$ induced by any arbitrary neighborhood function, there always exists a map to $\mathcal{C}_{2}$ where
$\forall\sigma,\tau\in\mathcal{C}_{1}\text{ if }\sigma\subseteq\tau\text{ then }f(\sigma)\subseteq f(\tau)$
holds trivially by selecting either an adjacency or an incidence since these neighborhood functions preserve set-composition and
$\forall\sigma\in\mathcal{C}_{1}$ (11)
holds since $f$ preserves the rank. However, by our initial construction $CC_{1}\not\simeq CC_{2}$ and thus we reach a contradiction. ∎
Appendix C Counting neighborhoods and message passing routes
@@
C.1 Neighborhood types in combinatorial complexes
Combinatorial representations allow for a variety of neighborhood functions, depending on how cells of different ranks interact. For a 2-dimensional combinatorial complex (with 0-, 11-, and $2$-cells), Figure˜3 illustrates all possible neighborhood functions. These relations can be grouped into two main groups:
@@ C.1 Neighborhood types in combinatorial complexes
C.2 Counting neighborhoods
The combinatorial complex depicted in Figure˜3 has rank $R=\dim({\mathcal{T}})=2$, which results in 12 distinct neighborhood functions. These neighborhood functions can be divided into the two categories from above: (i) incidence relations and (ii) adjacency relations.
@@ C.2 Counting neighborhoods
distinct neighborhood functions. For the example in Figure˜3, where $R=2$, this formula recovers the 12 neighborhoods observed.
C.3 Message passing routes
The number of neighborhood functions increases quadratically with the maximum rank, which may initially seem manageable, especially since most practical datasets involve low-rank structures (e.g., $R=2$ or $R=3$). However, the true combinatorial complexity arises when considering how the results neighborhoods can be composed. To formalize this notion, the following definition is introduced.
Definition C.1 (Message passing route).
A message passing route is an ordered set of neighborhood functions that a model uses to propagate information across a data structure.
Information may flow intra-rank, through adjacency relations, or inter-rank, through incidence relations. For example, a network might use the message passing route ${\mathcal{I}}_{2\to 1},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{0\to 2},{\mathcal{A}}_{0,1}]$ to transmit information from faces to edges, then from edges to nodes, from nodes back to faces, and finally between nodes connected by an edge.
@@ C.3 Message passing routes
Definition C.2 (Minimal message passing route).
@@
Minimal routes are of particular interest, as they guarantee full interdependence between all cells—a desirable property for expressive models. Moreover, minimal length implies simplicity in design, yielding the most compact routes that preserve this full dependency. For instance, the previously introduced route ${\mathcal{I}}_{2\to 1},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{0\to 2},{\mathcal{A}}_{0,1}]$ can be transformed into a minimal message passing route by removing the adjacency ${\mathcal{A}}_{0,1}$. In this reduced route, information from all cells becomes accessible to all others within two message passing steps. In fact, consider the case of 0-cells: after the first step, they depend on $1$-cells; after the second step, since $1$-cells now incorporate information from $2$-cells, the 0-cells become indirectly connected to them as well.
By definition, minimal message passing routes rely exclusively on incidence relations, as adjacencies do not propagate information inter-rank. Since each incidence targets a specific rank, and every rank must be updated once, minimal routes necessarily have length $R+1$. To enumerate all possible minimal routes, consider the set of ranks $T=\{0,1,\dots,R\}$, of size $R+1$. There are $(R+1)!$ distinct permutations of these ranks, each defining a unique order in which the ranks are updated. Given such a permutation $\{r_{0},r_{1},\dots,r_{R}\}$, a corresponding minimal message passing route can be constructed by selecting the sequence of incidence relations ${\mathcal{I}}_{r_{j-1}\to r_{j}}$ for $j=1,\dots,R$, and closing the cycle with ${\mathcal{I}}_{r_{R}\to r_{0}}$. This construction yields exactly one valid minimal route per permutation, resulting in a total of $N=(R+1)!$ such routes. In the case of graphs, where $R=1$, this results in exactly two minimal message passing routes.
C.4 Beyond minimal routes: including adjacencies
Minimal message passing routes represent only a small subset of the possible routes that can be constructed using the full set of neighborhoods in a combinatorial complex. A natural extension involves augmenting these minimal routes with adjacency operations, a common practice in many TDL architectures. For simplicity, the analysis here is restricted to the case where adjacencies are applied strictly after all incidence relations.